Tommy is 1/4 the age of his grandmother.Their ages total 75 years
How old is his grandmother?
How old would Tommy be in five years time?

Answer by richwmiller(17219)

4t=g
t+g=75
t+4t=75
5t=75
t=15
g=60
tommy will be 20=t+5 and gramma will be 65